# Sub-shot noise sensitivity via deformed four-headed kitten states

### v1

@@ Abstract
The original compass state, created by superposing four coherent states, yields anisotropic sub-Planck structures and demonstrates enhanced sensitivity to perturbations, offering advantages for quantum sensing. We propose two variants of this compass state by simultaneously applying photon addition and subtraction in different orders: one with addition first and one with subtraction first to the state. Our variants display sub-Planck structures and improved sensitivity to displacements, with photon addition and subtraction influencing these characteristics. In our cases, adding photons increases the average photon number, while photon subtraction lowers it in the first case and has no effect in the second. Furthermore, an increment in the added number of photons uniformly reduces the size of sub-Planck structures, whereas increasing the number of photons subtracted from the state causes these sub-Planck structures to expand in size; higher photon addition improves sensitivity, while photon subtraction decreases it. Remarkably, under optimal parameters, our specific variants achieve isotropic sub-Planck structures and provide isotropic enhanced sensitivity across all directions, surpassing compass states.
@@ I Introduction
Coherent states were first introduced by Schrödinger in 1926 [1, 2], and the concept was further developed in quantum optics by Glauber in 1963 [3]. Quantum superposition phenomena have been intensively investigated within the framework of the harmonic oscillator, leading to the creation of intriguing quantum states [4, 5]. Nonclassical nature of such states is revealed through nonclassical phase-space features, often visualized using the Wigner function [6, 7], and these states are considered valuable resources for continuous variable (CV) quantum information processing [8, 9, 10, 11, 12]. Notably, the macroscopic cat state [13, 14], which is the superposition of two distinct coherent states, is a prominent example of a nonclassical quantum state. The notion has been evolved to a generalized form of macroscopic cat states [15, 16, 17, 11]. A noteworthy example is the compass state [17], a concept that was introduced by Zurek in 2001. This state exhibits fascinating nonclassical phase-space features at scales significantly smaller than the Planck scale [18, 19]. These sub-Planck scale structures are critically significant due to their capacity to substantially increase sensitivity against displacement [20, 21]. The remarkable enhancement in displacement sensitivity afforded by compass states underscore their potential for advancing quantum measurement techniques and improving the performance of quantum technologies [22].
Both theoretical [23, 24, 25, 26] and experimental [27, 28, 29, 30] approaches have been employed to generate catlike states. Adding or subtracting photons is the key process by which matter interacts with light, allowing the creation of unique nonclassical states of light in experiments [31]. It has been demonstrated that applying photon addition and/or subtraction to squeezed-vacuum states may yield quantum states having phase-space features resembling those of a cat state, thereby providing an alternative approach for generating catlike states [32, 33, 34, 35]. Photon addition and subtraction have been performed simultaneously in different sequences on a coherent state [36, 37, 38]. These deformed coherent states may hold nonclassical and non-Gaussian phase-space characteristics that distinguish them from ordinary coherent states.
Compass states have been explored in different contexts [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 16, 51], and our analysis primarily focuses on the Zurek compass state [17]. This state is notable for their anisotropic sub-Planck phase-space structures and exhibits anisotropic enhanced sensitivity to displacements. In this work, we present variants of this compass state by applying photon addition and subtraction operations simultaneously in different sequences, resulting in various cases. Alternatively, our proposed quantum states can also be considered as a superposition of four deformed coherent states, exhibiting phase-space features similar to those of the original compass state. Our investigation employs a comprehensive phase-space formalism [7], incorporating Wigner function analysis and photon number distribution. Additionally, we also examine the sensitivity to displacements for each case [52, 53] and provide a detailed discussion of the physical significance and implications associated with each scenario. By optimizing the associated parameters, a number of our proposed cases may achieve improved phase-space characteristics relative to the compass state; that is, these cases contain isotropic versions of the sub-Planck structures, and the sensitivity enhancement is also isotropic.
@@
Our analysis is structured as follows: §II covers the basics of sub-Planck structures and their sensitivity to displacement, illustrated with the Zurek compass state. §III discusses deformed coherent states and our proposed quantum states, offering a detailed analysis. §IV discusses the physical basis and implications of our findings and provides a summary of our discussion.
@@
II Sensing of displacements
The quantum uncertainty principle for position and momentum, expressed through the commutation relation $[\hat{x},\hat{p}]:=\text{i}\hbar$ with $\hat{x}$ and $\hat{p}$ being the position and momentum operators, respectively, imposes constraints on the phase-space structure. Specifically, it dictates that the product of the uncertainties in position ($\Delta x$) and momentum ($\Delta p$) must satisfy the inequality $\Delta x\Delta p\geq\nicefrac{{\hbar}}{{2}}$ [18, 19]. The coherent state is known as the minimal uncertainty state [54], but a quantum state may also exhibit phase-space features that fall below the uncertainty limit and named as sub-Planck scale structures. The sensitivity to phase-space displacement of a quantum state may also connected to the smallest scale of its phase-space structure [21]. For example, the Zurek compass state [17] holds sub-Planck structures in phase space, which enhance its sensitivity beyond standard limits. This implies that the metrological capacity of a quantum state in certain situations may also link to its smallest phase-space structure. This connection will be examined in detail in this section, with a particular focus on the Zurek compass state as a primary example.
@@ II.1 Phase Space and Sensitivity
A Schrödinger coherent state can be expressed as a displaced vacuum state |$\ket{\alpha}:=\hat{D}(\alpha)\ket{0}$ with $\hat{D}(\alpha):=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a})$ is the displacement operator [55], where $\alpha\in\mathbb{C}$, and $\hat{a}$ ($\hat{a}^{\dagger}$) are annihilation (creation) operators. Coherent states are the most classical pure states of light [54], but their superposition may exhibit nonclassical aspects due to quantum interference [56]. The Wigner function denoted by $W_{\hat{\rho}}(\bm{\beta})$ with $\bm{\beta}:=(x,p)^{\top}$ constitutes phase space of a quantum state $\hat{\rho}$ [7], where $x$ and $p$ are the position and momentum pairs, respectively.
@@
A cat state is visible when the parameter $\alpha$ is high enough to maintain the orthogonality criterion $\braket{\alpha}{-\alpha}\approx 0$. If the constituent coherent states in this superposition are not distinguishable, the resulting states could be called Schrödinger-kitten states [34]. The compass state [17] is an extended and generalized form of this superposition, and can be considered a type of multi-component cat state [15, 59, 60, 61], and may also appear in different forms [16, 47, 48, 49].
Let us now discuss the concept of sensitivity to phase-space displacement. For a pure quantum state |$\ket{\psi}$, this sensitivity can be mathematically determined by evaluating the overlap function between a quantum state and its slightly displaced version. This involves calculating how much one state resembles another when displaced in phase space, which provides insight into how precisely the quantum state can detect or respond to changes in its phase-space configuration. Mathematically, this sensitivity can be determined by using [52]
@@
The phase-space structure and sensitivity to displacement of coherent states are at the standard limits, also known as shot noise limit [20]. The Wigner function of a macroscopic cat state typically manifests as two distinct Gaussian peaks in phase space with an oscillatory interference pattern, where each peak corresponds to a coherent state; this configuration may also termed as a two-headed cat state [62, 61]. In contrast, for a compass state, the Wigner function now hold four Gaussian peaks, catlike interference pattern and a super oscillatory pattern that reveals sub-Planck structures [17], may also classified as a four-component cat state [61, 60], aligning with the concept of multichotomous cat states [15]. Cat states, such as the one presented in Eq. (2), do not exhibit sub-Planck features as their interference phase-space attributes are not limited in all directions of phase space [47]. Different flavors of the compass states, as introduced in [16, 47, 48, 49], hold sub-Planck structures. Their sensitivity against displacement is also enhanced compared to the standard limits, thereby retaining sub-shot noise sensitivity [20]. We review these concepts with details next.
@@
II.2 Near-isotropic Sub-Planckness
Nonclassical effects in the Wigner function of a quantum system can give rise to remarkable phase-space structures, and these distinctive features play a significant role in enhancing sensitivity to phase-space displacements, a useful concept have been employed in quantum measurements [21, 20]. We now explore this concept through specific examples.
As an example, let us first examine the Wigner function of a coherent state, that is,
@@ II.2 Near-isotropic Sub-Planckness
Let us now include the example of the Zurek compass state [17], which is recognized as a superposition of coherent states given by $\alpha_{1}=\nicefrac{{c_{0}}}{{\sqrt{2}}}$, $\alpha_{2}=\nicefrac{{-c_{0}}}{{\sqrt{2}}}$, $\alpha_{3}=\nicefrac{{\text{i}c_{0}}}{{\sqrt{2}}}$, and $\alpha_{4}=\nicefrac{{-\text{i}c_{0}}}{{\sqrt{2}}}$, with $c_{0}\in\mathbb{R}^{+}$. This superposition can also be interpreted as the superposition of two cat states, or equivalently as the superposition of four coherent states and is denoted as
@@
We present $W_{\ket{\diamondsuit}}(\beta)$ in Fig. 1 for a few $c_{0}$ values. As depicted in Fig. 1(a), for $c_{0}=1$, the four coherent states in the compass state cannot be distinguished individually; hence, the corresponding Wigner distribution shows a central positive peak. This scenario represents the four-component kitten state, or alternatively may also referred as a four-headed kitten state [61]. However, for $c_{0}=5$, these four coherent states are now well separated and appear as four Gaussian lobes in the phase space, and the interference pattern is now pronounced in the phase space, as shown in Fig. (1)(b). Figure 1(c) shows that increasing the macroscopic parameter to $c_{0}=8$ causes coherent states to be pushed further away from the phase-space origin, resulting in enhanced negative regions in the intensity plot. Figure 2 illustrates the central interference pattern of each case presented in Fig. 1, where the phase-space features are arranged in a tiled format. Each tile in this pattern have an extension considerably smaller than that of a coherent state, indicating that these structures are at the sub-Planck scale. Furthermore, as $c_{0}$ increases, the size of the sub-Planck features decreases. These tile shape sub-Planck structures highlights their anisotropic properties, contrasting with an isotropic configuration where uniformity in all directions would be expected.
The enhancement in sensitivity for a cat state is limited to a particular direction in the phase space. Consequently, compass states are deemed more advantageous than cat states as they may offer greater sensitivity [47]. The overlap between the compass state $\ket{\diamondsuit}$ and its displaced version $\hat{D}(\delta)\ket{\diamondsuit}$ reflects the sensitivity to displacement in phase space
@@
In summary, considering the compass state under consideration, sensitivity to displacement appears to be associated to the macroscopic parameter $c_{0}$; as $c_{0}$ grows, it increases both sensitivity and the average photon number in the state. This suggests that a compass state with a larger average photon number can have greater susceptibility to displacement. Furthermore, in this example, the sensitivity to displacement is anisotropically amplified, as illustrated by the tile-like structures around the origin in Fig. 3.
III Photon-varied quantum states
Techniques involving photon addition (or subtraction) to squeezed-vacuum states have been effectively utilized to generate Schrödinger cat states [32, 33, 34, 35] and can also produce multi-component cat states [30, 49], highlighting the benefits of multi-photon operations. In our current work, we use photon addition and subtraction operations on quantum states to create innovative nonclassical quantum states. These newly proposed states could potentially align with modern experimental setups, as discussed in [34, 35]. In the following sections, we present the quantum states that are the focus of our study, accompanied by a thorough theoretical investigation. This detailed analysis explores the phase-space characteristics and implications of these quantum states, providing a comprehensive understanding of their characteristics and significance.
Figure 4: (a)-(c) represents Wigner distributions of the SA case of a coherent state, and (d)-(f) represents the corresponding AS cases: (a) $p=0$, $q=0$; (b) $p=4$, $q=2$; (c) $p=4$, $q=4$; (d) $p=0$, $q=2$; (e) $p=4$, $q=2$; and (f) $p=4$, $q=4$.
@@ III.1 Deformed Coherent States
Here, we examine the foundational concepts leading to our primary quantum states, specifically deformed coherent states [36, 37, 38]. These deformed coherent states are generated by applying a sequence of photon addition and subtraction, or vice versa, to a standard coherent state. We now explore how these operations modify the coherent state, resulting in the creation of deformed coherent states and their significance in our study.
The sequence of photon operations is crucial in our case, as different orders produce different results. Consider the scenario where $p$ photons are first added to a state followed by the subtraction of $q$ photons. This sequence appears as subtraction and addition (SA) operations, and is denoted appropriately for a coherent state as
|$\ket{\oplus}:=N^{-\nicefrac{{1}}{{2}}}_{\oplus}\hat{a}^{q}\hat{a}^{\dagger p}\ket{\alpha},$ (12)
@@
$N_{\oplus}=$ $(-1)^{p+q}\sum^{p}_{n=0}\Gamma H_{p-n,q}\left[\text{i}\alpha,\text{i}\alpha^{*}\right]$
×$\times H_{p-n,q}\left[\text{i}\alpha^{*},\text{i}\alpha\right]~{}\text{with}~{}\Gamma:=\frac{(-1)^{n}(p!)^{2}}{n![(p-n)!]^{2}}$ (13)
represents the normalization coefficient, and $H_{x,y}$ denotes the bivariate Hermite polynomial.
In the addition-subtraction (AS) scenario, the process involves first subtracting $q$ photons from a state, followed with the addition of $p$ photons. This sequence of operations, for a coherent state, can be denoted as
|$\ket{\ominus}:=N^{-\nicefrac{{1}}{{2}}}_{\ominus}\hat{a}^{\dagger p}\hat{a}^{q}\ket{\alpha}$ (14)
@@
$N_{\ominus}=\sum^{p}_{n=0}(-1)^{n}\Gamma H_{p-n,q}\left[\alpha,\alpha^{*}\right]H_{p-n,q}\left[\alpha^{*},\alpha\right]$ (15)
denotes the normalization coefficient.
The Wigner functions of SA and AS cases are evaluated using Eq. (1). Compared to their counterparts of ordinary coherent states, these Wigner functions attain non-Gaussian form and may hold negative phase-space attributes for different values of $p$ and $q$, highlighting their nonclassical nature [36].
@@
$W_{\ket{\oplus}}(\beta)$ $=\frac{1}{N_{\oplus}}W_{\ket{\alpha}\bra{\alpha}}(\beta)\sum^{p}_{n=0}\Gamma H_{p-n,q}\left[-\text{i}\Omega,-\text{i}\alpha^{*}\right]$
×$\times H_{p-n,q}\left[\text{i}\Omega^{*},\text{i}\alpha\right]~{}\text{with}~{}\Omega:=2\beta-\alpha.$ (16)
The non-Gaussian nature of this Wigner function is obvious from the expression, and as it is also evident in Figs. 4(a)-(c) that this Wigner function now acquired negative amplitudes, suggesting the nonclassical nature of this state. Here, note that Fig. 4(a) with $p=0$ and $q=0$, presents the case of corresponding ordinary coherent state, but as observed in Figs. 4(b) and 4(c), when $p$ grows, negative regions are amplified, while the $q$ parameter has the opposite impact.
Mathematically, the Wigner function for the AS case has an analogous form to that of the SA case,
$W_{\ket{\ominus}}(\beta)=$ $\frac{1}{N_{\ominus}}W_{\ket{\alpha}\bra{\alpha}}(\beta)\sum^{p}_{n=0}\Gamma H_{p-n,q}\left[\Omega^{*},\alpha\right]$
×$\times H_{p-n,q}\left[\Omega,\alpha^{*}\right].$ (17)
This Wigner function is shown in Figs. 4(d)-(f) with different $p$ and $q$ values, indicating that the corresponding Wigner functions also contains negative amplitudes in the phase space. It is interesting to note that when $q$ is applied directly on a coherent state, there is no effect, as shown in Fig. (4)(d) with $p=0$ and $q=2$ appearing to have the same Wigner function as a coherent state. As illustrated in Fig. (4)(e), similar to SA case, the photon addition in this case also enhances the negative regions. However, as can be observed in Fig. (4)(f), with a nonzero value of $p$, the parameter $q$ appears to be effective and serves the same as in the previous SA example, that is, it appears to make the negative regions less dense. The effects of $p$ and $q$ on the phase space are more thoroughly discussed for the quantum states presented in the following sections.

### v2

@@ Abstract
We explore nonclassical effects in the phase space of a four-headed kitten state (a superposition of two Schrödinger kitten states) induced by photon addition and subtraction operations applied in different sequences. We investigate two scenarios: in the first, photon addition is applied to the state, followed by photon subtraction, while in the second, the order of operations is reversed. We demonstrate that applying multiphoton operations to the state results in notable nearly isotropic sub-Planck structures, with the characteristics of these structures being influenced by the photon addition and subtraction. We observe that adding photons increases the average photon number, while photon subtraction reduces it in the first case but has no effect in the second. Increasing the number of added photons compresses the sub-Planck structures in both cases. Photon subtraction, however, has the opposite effect on the sub-Planck structures in the first case and no effect in the second, although it may improve their isotropy at optimal settings. The presence of the sub-Planck structures in our states leads to improved sensitivity to displacements, exceeding the standard quantum limit, as verified across all the depicted scenarios.
@@ I Introduction
Coherent states were first introduced by Schrödinger in 1926 [1, 2], and the concept was further developed in quantum optics by Glauber in 1963 [3]. Quantum superposition phenomena have been intensively investigated within the framework of the harmonic oscillator, leading to the creation of intriguing nonclassical states [4, 5]. The nonclassical nature of quantum states is revealed through nonclassical phase-space features, often visualized using the Wigner function [6, 7]. Nonclassical states are considered valuable resources for continuous variable quantum information processing, particularly for sensing and metrology applications [8, 9, 10, 11, 12]. A critical factor in determining the utility of a quantum state for such applications is the degree to which it becomes distinguishable from the initial state after a small displacement, which is influenced by the smallest phase-space feature of a quantum state and measures sensitivity to displacement (perturbation) [13, 14, 15]. Developing quantum states with finer phase-space features is essential for enhancing their capacity to measure smaller scale displacements, and quantum states with finer phase-space features are highly valuable for advancing quantum measurement techniques and improving the performance of quantum technologies [16].
The phase-space volume occupied by a coherent state adheres to the Planck scale $\hbar$ [17], while also preserving its sensitivity to displacements at this standard limit [18, 19]. It has been shown that specific chaotic quantum systems [13, 20] are able to produce phase-space features with volumes far below this standard norm. A macroscopic cat state [21, 22], which is the superposition of two distinct coherent states, is a prominent example of nonclassical states and has shown more finer phase-space features and greater sensitivity compared to a coherent state [23]. The notion of coherent-state superpositions has been evolved to a generalized form of macroscopic cat states [24, 25, 13, 11]. Compass states [26, 27, 28, 29, 30, 31, 32, 33, 34, 23, 35, 36, 37, 25, 38, 13] (a superposition of four coherent states) have been shown to exhibit improved phase-space characteristics compared to their precursors, namely coherent states and cat states. In particular, it has been found that these states hold sub-Planck scale features (dimension below than the Planck scale) and hold sensitivity to displacements greater than the standard quantum limit, which make them a potential candidate for quantum metrology applications [39]. Both theoretical [40, 41, 42, 43, 44] and experimental [45, 46, 47, 48] approaches have been employed to generate catlike states.
Multiphoton operations applied to a quantum state provide innovative techniques for controlling fundamental nonclassical characteristics [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. For example, it has been shown that photon subtraction (or addition) from ordinary squeezed-vacuum states may develop quantum states similar to those of cat states [50, 53, 58, 52]. In this work, we consider a four-headed kitten state, which is the superposition of two kitten states deficient in nonclassical phase-space attributes. Photon addition and subtraction operations are applied on this state in different orders to build new quantum states, which exhibit intriguing phase-space characteristics similar to a compass state [13]. In our first case, we apply photon addition followed by photon subtraction to the state, while in the second case, the order of operations is reversed, with photon subtraction applied first and then photon addition.
Our investigation utilizes phase-space formalism [7], incorporating Wigner function analysis and photon number distributions. We compare the states to explore the effects of multiphoton operations and the optimal selection of parameters, followed by an analysis of their sensitivity to displacements. The analysis concludes with a thorough discussion of the physical significance and implications of each scenario. Specifically, we show that multiphoton operations applied on the four-headed kitten state lead to quantum states with refined phase-space characteristics, and interestingly, our proposed cases also hold sub-Planck phase-space features whose occurence is now affilated with the number of added and subtracted photons. Furthermore, we show that our proposed instances hold sensitivity to phase-space displacements greater than the standard quantum limit, making them a promising choice for quantum sensing applications. Our observations indicate that as the number of added photons increases, finer sub-Planck structures develop in our cases. Conversely, photon subtraction in the first scenario causes the sub-Planck structures to expand. In the second scenario (subtraction followed by addition), photon subtraction becomes ineffective, but in this case, an improvement in the isotropy (directional invariance in phase space) of the sub-Planck structures is observed for certain photon subtraction choices. These findings align with the behavior observed in the sensitivity of our quantum states.
@@
Our paper is structured as follows: §II reviews the basics of sub-Planck structures and their impact on sensitivity to displacement, illustrated with the Zurek compass state. §III discusses deformed coherent states and our proposed quantum states, offering a detailed analysis to identify the corresponding phase-space characteristics. §IV presents phase-space sensitivity contained by our proposed quantum states. §V provides a compact summary of our findings, and §VI provides the main conclusions of our results.
@@
II Concepts and perspectives
The quantum uncertainty principle for position and momentum, expressed through the commutation relation $[\hat{x},\hat{p}]:=\text{i}\hbar$ with $\hat{x}$ and $\hat{p}$ being the position and momentum operators, respectively, imposes constraints on the dimension of a phase-space structure. Specifically, it dictates that the product of the uncertainties in position ($\Delta x$) and momentum ($\Delta p$) satisfies $\Delta x\Delta p\geq\nicefrac{{\hbar}}{{2}}$ [18, 19]. This implies that the dimension of a phase-space feature is constrained by this standard limit; if a phase-space feature has dimensions below this threshold, it is considered unphysical. This general assumption is challenged by Zurek [13], who demonstrates that the sub-Planck structures in the compass state significantly influence the phase-space properties of the states. These spotty features have been found crucial in enhancing sensitivity to displacements, and this improvement in sensitivity is directly dedicated to the sub-Planck scale structures in the state. In this section, we primarily focus on reviewing the key concept of sub-Planck structures and their critical role in enhancing sensitivity to displacements.
@@ II.1 Phase space and sensitivity
A Schrödinger coherent state can be expressed as a displaced vacuum state |$\ket{\alpha}:=\hat{D}(\alpha)\ket{0}$ with $\hat{D}(\alpha):=\exp(\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a})$ the displacement operator [61], where $\alpha\in\mathbb{C}$, and $\hat{a}$ ($\hat{a}^{\dagger}$) are annihilation (creation) operators. Coherent states are the most classical pure states of light [17], but their superposition may exhibit nonclassical aspects due to quantum interference [62]. The Wigner function denoted by $W_{\hat{\rho}}(\bm{\beta})$ with $\bm{\beta}:=(x,p)^{\top}$ constitutes phase space of a quantum state $\hat{\rho}$ [7], where $x$ and $p$ are the position and momentum pairs, respectively.
@@
A cat state is visible when the parameter $\alpha$ is high enough to maintain the orthogonality criterion $\braket{\alpha}{-\alpha}\approx 0$. If the constituent coherent states in this superposition are not distinguishable, the resulting states is known as Schrödinger-kitten states [52].
We now discuss the concept of sensitivity to phase-space displacement. For a pure quantum state |$\ket{\psi}$, this sensitivity can be mathematically determined by evaluating the overlap function between a quantum state and its slightly displaced version. This involves calculating how much one state resembles another when displaced in phase space, which provides insight into how precisely the quantum state can detect or respond to changes in its phase-space configuration. Mathematically, this sensitivity can be determined by using [14]
@@
The Wigner function of a macroscopic cat state typically manifests as two distinct Gaussian peaks in phase space with an oscillatory interference pattern, where each peak corresponds to a coherent state; this quantum state is simply called a macroscopic cat state [65, 66]. In contrast, for a compass state, the Wigner function now holds four Gaussian peaks, a catlike interference pattern, and central sub-Planck structures [13], may also be classified as a four-component cat state [66, 67], aligning with the concept of multichotomous cat states [24]. Cat states, such as the one presented in Eq. (2), do not exhibit sub-Planck features as their interference phase-space features are not limited in all directions of phase space [23].
@@
II.2 Sub-Planckian sensitivity
The Wigner function of a coherent state is denoted as
@@ II.2 Sub-Planckian sensitivity
We now include the example of the Zurek compass state [13], which is recognized as a superposition of coherent states given by $\alpha_{1}=\nicefrac{{c_{0}}}{{\sqrt{2}}}$, $\alpha_{2}=\nicefrac{{-c_{0}}}{{\sqrt{2}}}$, $\alpha_{3}=\nicefrac{{\text{i}c_{0}}}{{\sqrt{2}}}$, and $\alpha_{4}=\nicefrac{{-\text{i}c_{0}}}{{\sqrt{2}}}$, with $c_{0}\in\mathbb{R}^{+}$. This superposition can also be interpreted as the superposition of two cat states, or equivalently as the superposition of four coherent states and is denoted as
@@
We present $W_{\ket{\diamondsuit}}(\beta)$ in Fig. 1 for a few $c_{0}$ values. As depicted in Fig. 1(a), for $c_{0}=1$, the four coherent states in the compass state cannot be distinguished individually; hence, the corresponding Wigner distribution shows a central positive peak. This scenario represents the four-component kitten state, or alternatively may also referred as a four-headed kitten state [66]. However, for $c_{0}=5$, these four coherent states are now well separated and appear as four Gaussian lobes in the phase space, and the interference pattern is now pronounced in the phase space, as shown in Fig. (1)(b). Figure 1(c) shows that increasing the macroscopic parameter to $c_{0}=8$ causes coherent states to be pushed further away from the phase-space origin, resulting in enhanced negative regions in the intensity plot. Figure 2 illustrates the central interference pattern of each case presented in Fig. 1, where the phase-space features are arranged in a tiled format. Each tile in this pattern have an extension considerably smaller than that of a coherent state, indicating that these structures are at the sub-Planck scale. Furthermore, as $c_{0}$ increases, the size of the sub-Planck features decreases.
The enhancement in sensitivity for a cat state is limited to a particular direction in the phase space. Consequently, compass states are deemed more advantageous than cat states as they may offer greater sensitivity [23]. The overlap between the compass state $\ket{\diamondsuit}$ and its displaced version $\hat{D}(\delta)\ket{\diamondsuit}$ reflects the sensitivity to displacement in phase space,
@@
In summary, for the compass state under consideration, sensitivity to displacement appears to be associated with the macroscopic parameter $c_{0}$; as $c_{0}$ grows, it increases both sensitivity and the average photon number in the state. This suggests that a compass state with a larger average photon number can have greater susceptibility to displacement. Furthermore, in this example, the sensitivity to displacement is anisotropically amplified, as illustrated by the tile-like structures around the origin in Fig. 3. Our main focus is on the four-headed kitten state, which is shown in Fig. 1(a). It is evidenced that this particular case is devoid of negative amplitudes, and the dimensions of its phase-space feature are also comparable to those of the coherent states given that sub-Planck structures are absent in this case, and our multiphoton illustrations are specifically devoted to this case, and we will demonstrate how these photon operations will alter its phase-space characteristics by keeping $c_{0}$ a constant.
Figure 4: The overlap between the coherent and deformed versions: (a) For the SA version of the coherent state, the red dotted line represents $r=4$, $q=0$, and the blue solid line represents $r=4$, $q=4$; (b) For the AS version of the coherent state, the red dotted line corresponds to $r=4$, $q=0$, and the blue dotted line represents $r=4$, $q=4$. The horizontal black solid lines indicate the case where $r=0$ and $q=0$, representing the overlap between two ordinary coherent states.
III Refined phase-space densities
Multiphoton operations have been used both theoritically and experimentally to enhance the phase-space characteristics of a quantum state. Techniques involving photon addition (or subtraction) to squeezed-vacuum states have been effectively utilized to generate Schrödinger cat states [50, 51, 52, 53] and can also produce multi-component cat states [48, 36], highlighting the benefits of multi-photon operations. In our current work, we use photon addition and subtraction operations on quantum states to create innovative nonclassical quantum states. These newly proposed states could potentially align with modern experimental setups, as discussed in [52, 53]. In the following sections, we present the quantum states that are the focus of our study, accompanied by a thorough theoretical investigation. This detailed analysis explores the phase-space characteristics and implications of these quantum states, providing a comprehensive understanding of their characteristics and significance.
Figure 5: (a)-(c) represents Wigner distributions of the SA case of a coherent state, and (d)-(f) represents the corresponding AS cases: (a) $r=0$, $q=0$; (b) $r=4$, $q=2$; (c) $r=4$, $q=4$; (d) $r=0$, $q=1$; (e) $r=4$, $q=2$; and (f) $r=4$, $q=4$. In all cases $\alpha=\nicefrac{{1}}{{\sqrt{2}}}$.
@@ III.1 Deformed coherent states
In this section, we review the basic concepts that lead to our primary quantum states; especially, we discuss the concept of deformed coherent states [54, 55, 56]. Deformed coherent states can be generated by applying a sequence of photon addition and subtraction, or vice versa, to a standard coherent state.
Consider the scenario where $r$ photons are first added to a state, followed by the subtraction of $q$ photons. This sequence appears as subtraction and addition (SA) operations and is denoted for a coherent state as
|$\ket{\oplus}:=N^{-\nicefrac{{1}}{{2}}}_{\oplus}\hat{a}^{q}\hat{a}^{\dagger r}\ket{\alpha},$ (12)
@@
$N_{\oplus}=$ $(-1)^{r+q}\sum^{r}_{n=0}\Gamma H_{r-n,q}\left[\text{i}\alpha,\text{i}\alpha^{*}\right]$
×$\times H_{r-n,q}\left[\text{i}\alpha^{*},\text{i}\alpha\right]~{}\text{with}~{}\Gamma:=\frac{(-1)^{n}(r!)^{2}}{n![(r-n)!]^{2}}$ (13)
represents the normalization coefficient and $H_{x,y}$ denotes the bivariate Hermite polynomial.
In the addition-subtraction (AS) scenario, the process involves first subtracting $q$ photons from a state, followed with the addition of $r$ photons. This sequence of operations for a coherent state can be denoted as
|$\ket{\ominus}:=N^{-\nicefrac{{1}}{{2}}}_{\ominus}\hat{a}^{\dagger r}\hat{a}^{q}\ket{\alpha}$ (14)
@@
$N_{\ominus}=|\alpha|^{2q}\sum^{r}_{n=0}(-1)^{n}\Gamma|\alpha|^{2(r-n)}$ (15)
denotes the corresponding normalization coefficient.
To illustrate the impact of photon addition and subtraction operations on a coherent state, we calculate the overlap between the coherent state and its SA and AS variants. The overlap between the coherent state |$\ket{\alpha}$ and its SA version $\ket{\oplus}$ is denoted as $F_{\oplus}(\alpha):=|\braket{\alpha}{\oplus}|^{2}$, while the overlap in the AS case is denoted as $F_{\ominus}(\alpha):=|\braket{\alpha}{\ominus}|^{2}$. These overlaps are evaluated numerically and are represented in Fig. 4. This illustrates that both SA and AS situations differ significantly from the coherent in lower $\alpha$ ranges, and as $\alpha$ grows, corresponding SA and AS variants of the coherent state simply transofrm back to the ordinary coherent state. This comparison is carried out with different amounts of photon addition $r$ and subtraction $q$, as shown in Fig. 4(a) and Fig. 4(b), respectively. Note that the photon subtraction operation in the AS scenario has no effect, and in this case, it essentially represents the photon-added version of the coherent state. This is shown in Fig. 4(b), where the red dotted line and blue dotted line indicate different $q$ values with the same $r$ values, resulting in the same curve.
The Wigner functions of SA and AS cases can be evaluated using Eq. (1). Compared to their counterparts of ordinary coherent states, these Wigner functions attain non-Gaussian form and may hold negative phase-space attributes for different values of $r$ and $q$, highlighting their nonclassical nature [54].
@@
$W_{\ket{\oplus}}(\beta)=$ $\frac{1}{N_{\oplus}}W_{\ket{\alpha}}(\beta)\sum^{r}_{n=0}\Gamma H_{r-n,q}\left[-\text{i}\Omega,-\text{i}\alpha^{*}\right]$
×$\times H_{r-n,q}\left[\text{i}\Omega^{*},\text{i}\alpha\right],$ (16)
where
$\Omega=2\beta-\alpha.$ (17)
The non-Gaussian nature of this Wigner function is obvious from the expression, and as it is also evident in Figs. 5(a)-(c) that this Wigner function now acquired negative amplitudes, suggesting the nonclassical nature of this state. Here, note that Fig. 5(a) with $r=0$ and $q=0$, presents the case of corresponding ordinary coherent state, but as observed in Figs. 5(b) and 5(c), when $r$ grows, negative regions are amplified, while the $q$ parameter has the opposite impact. Mathematically, the Wigner function for the AS case has an analogous form to that of the SA case.
$W_{\ket{\ominus}}(\beta)=$ $\frac{(-1)^{r}}{N_{\ominus}}W_{\ket{\alpha}}(\beta)\sum^{r}_{n=0}(-1)^{n}\Gamma(2\beta-\alpha)^{r-l}$
×$\times(\alpha^{*}-2\beta^{*})^{r-l}.$ (18)
This Wigner function is shown in Figs. 5(d)-(f) with different $r$ and $q$ values, indicating that the corresponding Wigner functions also contain negative amplitudes in the phase space. Note that in this case $q$ has no effect on the state, as shown in Fig. (5)(d) with $r=0$ and $q=2$ appearing to have the same Wigner function as a coherent state. This is additional evidence that the photon subtraction effect in the indicated AS case is zero, which simply reflects the photon-added case. The photon subtraction from a coherent state leaves the state uncnhaged is experimentally proved [57]. As illustrated in Fig. (5)(e), similar to the SA case, the photon addition in this case also enhances the negative regions. Furthermore, as observed in Fig. (5)(f), with a nonzero value of $r$, the parameter $q$ remains ineffective. The effects of $r$ and $q$ on the phase space are more thoroughly discussed for the quantum states presented in the following sections.